Jack is 20 years older than Tom. 4 years later, Jack will be 3 times as old as Tom. How old is Jack now?
Answer by mananth(16946)

Tom's age x
Jack's age x + 20
4 years later
Tom's age x + 4
Jack's age x + 20 + 4 = x + 24
Jack is 3 times older
x+24=3(x+4)
x+24=3x +12
3x-x=24 -12
2x=12
/2
x= 6 Tom's age
x+20 =26 years Jack's age